OpenStudy (anonymous):

what is (5^4/5 * 5^4/5)^-10

OpenStudy (anonymous):

First, if there's a common base and they're being multiplied, just add the exponents and combine them. Then remove the negative on the 10 by putting it over one. Then distribute the exponent. \[(5^{4/5} \times 5^{4/5})^{-10} = (5^{8/5})^{-10}\] = \[1\over(5^{8/5})^{10}\] = \[1\over5^{80/5}\] = \[1\over5^{16}\]

OpenStudy (anonymous):

so you don' multiply the the 2 5's

OpenStudy (anonymous):

By multiplying the 2 fives, you add the exponents. It's a rule.\[a ^{b} \times a ^{c} = a ^{b + c}\]

OpenStudy (anonymous):

o well thanks a ton

OpenStudy (anonymous):

You're welcome. :)